OpenStudy (anonymous):

MEDAL AND FAN!! Write a two-step equation that needs the steps of addition then division in order to be solved, and then solve it. Show your work.

OpenStudy (anonymous):

\[20+4/2=x\] First you'd have to add 4 to 20 getting 24, then divide the 24 by 2. :)

OpenStudy (anonymous):

\[3x-5=25\]Need to add 5 then divide by 3 \[3x-5+5=25+5\]\[3x=30\]\[\frac{3x}{3}=\frac{30}{3}\]\[x=10\]